OpenStudy (anonymous):

Zaniya,Safee,and ken went shopping on Sunday. Together, they spent a total of $140. Zaniya spent 3 times as much as safee. Ken spent $5 more than Safee. How much did each person spend ?

OpenStudy (loser66):



OpenStudy (anonymous):

@loser66 I'm not able to see the picture properly

OpenStudy (anonymous):

Zaniya,Safee,and ken Safee spent $x Zaniya spent $3x Ken spent $x+5

OpenStudy (anonymous):

Total = 140

OpenStudy (anonymous):

Therefore, x+3x+x+5 = 140 4x+x+5 = 140 4x+x = 140 - 5 5x = 135 x = 135/5 x = 27 Now as, Safee spent $x Zaniya spent $3x Ken spent $x+5 Therefore, Safee spent $27 Zaniya spent 27*3 = $81 Ken spent = x+5 = 27 + 5 = $32 To verify,we do,this, 27 + 81 + 32 108+32 140 Hence,verified.

OpenStudy (anonymous):

@ChiefArnav thank you so much !!!!! :) i really appreciated that ! :)